OpenStudy (anonymous):

I need to get 400,000 (i wish) what would i have to deposit? if i was to get 6% compounded annually for 25 years what is the formula i would have to use? thanks in advance

OpenStudy (anonymous):

example in book is making no sence to me at all

OpenStudy (anonymous):

my mr jim ge sir helpful

jimthompson5910 (jim_thompson5910):

A = P(1+r/n)^(n*t) 400000 = P(1+0.06/1)^(1*25) 400000 = P(1+0.06)^(25) 400000 = P(1.06)^(25) 400000 = P(4.2918707) I'll let you finish up

OpenStudy (anonymous):

its the principle that for some reason i am confused on

jimthompson5910 (jim_thompson5910):

you would divide both sides by 4.2918707

jimthompson5910 (jim_thompson5910):

to solve 400000 = P(4.2918707) for P

OpenStudy (anonymous):

ty will try that

jimthompson5910 (jim_thompson5910):

yw